OpenStudy (anonymous):

The James family and the Jones family each used their sprinklers last summer. The water output rate for the James family's sprinkler was 25L/ per hour. The water output rate for the Jones family's sprinkler was 15L/ per hour. The families used their sprinklers for a combined total of 55 hours, resulting in a total water output of 1225L . How long was each sprinkler used?

OpenStudy (anonymous):

@jim_thompson5910 @amistre64 @mathmate @austinL

OpenStudy (austinl):

25J + 15L = 1225 J + L = 55 I believe this is what it is asking.

OpenStudy (ybarrap):

$$ \text{Let M = number of hours the James Family used their sprinkler}.\\ \text{Let N = number of hours the Jone's Family used their sprinkler.}\\ \text{The total volume used by both families is then }25M+15N=1225.\\ \text{They used the sprinklers a total of }M+N=55 \text{ hours.}\\ \text{Two equations, two unknowns.}\\ \text{Any questions?} $$